# Extended Resolution Clause Learning via Dual Implication Points

### v1

@@ 5 Extension Variables from Dual Implication Points
5.1 An example with a grid Tseitin principle
Given a graph where every vertex has a charge, a number which is 00 or $1$, a Tseitin formula [37] is created by considering one variable per edge, and adding one constraint per vertex $v$ expressing that the sum of the variables of all edges incident to $v$ modulo 2 is equal to its charge. The CNF version that we consider converts each xor constraint into clauses by simple enumeration of falsifying assignments. It is easy to see that the formula is unsatisfiable if and only if the sum of all charges is even. Here we consider as a graph the $3\times 3$-grid depicted in Figure 4 and only one vertex has charge 1, so the clauses are unsatisfiable.
@@ 5.1 An example with a grid Tseitin principle
Figure 4: An instance of the $3\times 3$-grid Tseitin principle. The nodes are assigned a polarity in $\{0,1\}$. The edges are labeled with variables.
@@
Figure 5: The complete conflict graph from the example; (a) at the first conflict and (b) at the second conflict. Boxed literals are decision literals. Implications at the top decision level are in bold.
It is interesting to relate the Figure 5(b) to Theorem 4.3 on DIPs. In this example, there is only one way to choose the two vertex-disjoint paths. Namely, to let $\pi_{a}$ and $\pi_{b}$ be the paths $e_{6},e_{8},e_{11},\overline{x},\overline{e_{7}},e_{10},\perp$ and $e_{6},e_{9},\overline{e_{12}},\perp$. The edge from $e_{6}$ to $\overline{e_{7}}$ bypasses $e_{8}$, $e_{11}$ and $\overline{x}$; and the edge from $e_{11}$ to $\overline{e_{12}}$ is a crossing separator. (The edge from $\overline{e_{7}}$ to $e_{9}$ is a “vacuous” crossing separator that does not actually remove any possible DIP pairs.)
@@
5.2 Extending CDCL with Dual Implication Points
@@
Filtering out bad-quality DIPs. Learning a DIP whenever we find one would be too prolific and overwhelm the solver. In order to determine whether the DIP chosen in the previous step had to be discarded, the first possibility we explored was again inspired by 1-UIP learning, where learning glue clauses, i.e. having LBD equal to $2$, is a desired situation. The first filtering mechanism we implemented discarded all DIPs that did not have a glue post-DIP clause. Another possibility we considered is to wait for a DIP to occur a certain number of times before using it in DIP-based learning. In our implementation, we tried $20$ and $5$ occurrences, as the minimum threshold to introduce a DIP. A third possibility is to use the activity-based heuristic of the literals in the DIP to assess its quality: DIPs whose literals have high decision-heuristic scores should be prioritized. In our implementation, we check whether the activity level of the current DIP is higher than the average activity level of the 20 most recently encountered DIPs; if so, the current DIP is a candidate for DIP-learning, otherwise it is discarded. We refer to these various techniques as glue, occ5 and act, respectively.
@@
It would make no sense to backtrack to level $\ell_{C}$ when $\ell_{C}<\ell_{D}$ since then neither $\lnot z$ nor $\lnot f$ would be propagated. However, another possible strategy would be to backtrack to the maximum of the decision levels $\ell_{C}$ and $\ell_{D}$. This would mean $\lnot z$ and $\lnot f$ are both propagated. The disadvantage of this when $\ell_{C}>\ell_{D}$ is that it would mean $\lnot z$ is asserted by the post-DIP clause at level $\ell_{C}$, whereas it could have been propagated at the previous decision level $\ell_{D}$. This breaks a usual invariant for CDCL solvers. This would only be possible in a solver that permits chronological backtracking [30, 29]; xMapleLCM, however, does not support this.
The previous reasoning needs some clarification for the case when the extension variable $z$ with definition $z\leftrightarrow l_{1}\land l_{2}$ we want to introduce already exists. If $z$ is undefined or defined at the current decision level nothing changes. We know it cannot be true at some previous level, because otherwise $l_{1}$ and $l_{2}$ would have been propagated at that level, and not in the current one as all literals belonging to a DIP do. If it is false at some previous level, then we have no guarantee that the pre or the post-DIP clause is asserting at any decision level. Fortunately, that rarely happens in practice. However, we can always perform standard 1UIP learning or try to apply DIP-based conflict analysis starting with the clause $z\lor\lnot l_{1}\lor\lnot l_{2}$ that we can guarantee is conflicting at the current decision level.
@@
Deleting extended variables. As it happens with lemma learning, where keeping too many lemmas slows down unit propagation, managing too many extended variables might also be counterproductive. Again following the analogy with learned lemmas, which are useful at some point of the search but might become inactive after a while, it is natural to think that extended variables follow the same behavior. All in all, it seems mandatory to consider the deletion of extended variables.
@@ 6 Experimental Evaluation
We have implemented the DIP-based clause learning schemes described in Section 5.2 on top of the xMapleLCM ER framework [^3] All sources used for this evaluation can be found in https://github.com/chjon/xMapleSAT/tree/main.. We started our experimental evaluation running a variant of DIP-based learning on all benchmarks of the 2023 SAT Competition [6] and comparing it to MapleLCM [25], the CDCL SAT solver on which it is based. Even though there did not seem to be a systematic improvement on all benchmarks, a few families with important speedups were identified. We start this section by analyzing the impact of DIP-based learning on these families and then move to final considerations about the performance on the overall 2023 SAT competition benchmarks.
6.1 Performance Analysis Methodology
For each family, we start by describing the problem they encode. After that, we report on the performance of a variety of state-of-the-art solvers, each with some distinguished characteristic: Kissat 3.1.1 [8] (an extremely efficient CDCL solver), CryptoMiniSat 5.11 [36] (support for XOR reasoning), SBVA+CaDiCaL [14] (introduction of new variables via SBVA and winner of the main track of the 2023 SAT Competition), GlucoseER [3] (extended-resolution based CDCL solver) and a variant of xMapleLCM-DIP, which we refer to as the baseline, that we describe next.
@@ 6.1 Performance Analysis Methodology
Finally, we evaluate the impact of the different techniques explained in Section 5.2. In particular, we first consider as a baseline a version that (i) finds DIPs in the middle of the conflict graph, (ii) only adds a DIP if it has occurred at least 20 times and (iii) always learns the both pre- and post-DIP learning clauses. Different variants can be obtained by changing only one of the three previous design decisions at a time. Regarding the type of DIP used, we analyze the performance of closest and random, that is, the systems whose only difference w.r.t. baseline is that the type of DIP is changed. Regarding the criterion used to discard a DIP, we analyze the glue, act and occ5 configurations, where the latter discards any DIP that has not occurred at least 5 times. Finally, the 1-clause variant only learns the post-DIP clause. Our goal is to understand why certain configurations do not work well on particular families.
6.1.1 Matching of Properly Intersecting Intervals
We are given a sequence of numbers $(a_{1},\ldots,a_{n})$, initially all set to zero, and a set of operations, each consisting of assigning a certain number to a contiguous subsequence of the $a_{i}$’s, defined by an interval. The goal is to perform each operation exactly once while maximizing the number of pairs of consecutive numbers $(a_{i},a_{i+1})$ that are different at the end of the process. Given some additional conditions, this can be formulated as maximum bipartite matching problem on a certain graph. By removing some edges from this graph, an unsatisfiable problem is generated. A more detailed description can be found in [6].
We downloaded the 23 instances submitted to the 2023 SAT Competition, using the Global Benchmark Database [20], and observed that no system without extended-resolution reasoning could solve any benchmark in a time limit of 1 hour. On the other hand, there are 14 benchmarks that GlucoseER could solve, of which our DIP-based variants could solve 7. Detailed results for our variants can be seen in Table 1.
Baseline 1-Clause Closest Rand Glue Occ5 Act
PROPERLY INTERSECTING INTERVALS FORMULAS
SOLVED 6 5 7 7 0 0 0
BEST 1 1 2 3 0 0 0
Median (secs.) 2777 2566 2358 2189 3600 3600 3600
TSEITIN FORMULAS
SOLVED 11 13 13 11 0 15 3
BEST 2 5 0 0 0 8 0
Median (secs.) 897 628 2707 2782 3600 425 3600
(X)ORIFIED RANDOM $k$-xor FORMULAS
SOLVED 10 8 11 9 0 9 7
BEST 5 2 1 2 0 1 0
Median (secs.) 430 803 1080 820 3600 936 3112
Table 1: Performance of DIP-learning variants on three sets of formulas. Median running times are in seconds and, for each row, the best performing system is in bold.
For this particular type of formulas, the rand variant performs the best, but detailed results indicate that we cannot claim it is systematically faster. One clear conclusion we can infer is that glue, occ5 and act exhibit very poor performance. A more detailed analysis on the runs of these systems reveal that, at least in these benchmarks, both act and occ5 apply DIP learning in many more conflicts than the other variants, whereas the distinctive characteristic of glue is that the percentage of decisions on extended variables was extremely low. As can be seen at the end of this section, the latter is a situation than tends to indicate the DIP learning is not helping the solver in this context.
6.1.2 Tseitin Formulas
These formulas have already been described in Section 5.1. In this section, we consider unsatisfiable formulas generated by CNFgen [24]. We started generating instances with grids of size $n\times n$ and ran all systems with a time limit of 300 seconds in order to analyze how different solvers scale. The performance of all systems can be seen on the left plot in the first row of Figure 6 (note the logarithmic scale on the $y$ axis). A point $(x,y)$ in a solver line indicates that the solver took $y$ seconds to solve the Tseitin formula on grid size $x\times x$.
@@ 6.1.2 Tseitin Formulas
Figure 6: Performance analysis of different solvers on Tseitin formulas.
These formulas are easy for CryptoMiniSat, since the application of Gaussian elimination in a preprocessing step solves them. Regarding the rest of the systems, only our DIP-based tool and GlucoseER have good scaling properties. We want to remark that this behavior is not unique to Tseitin formulas on the grid. We generated Tseitin formulas over random $4$-regular and $6$-regular graphs and the conclusions are the same as can be seen in the other two plots in the first row of Figure 6. These results start suggesting that, even though GlucoseER and our DIP-based learning are quite different approaches to integrating ER into CDCL solvers, they seem to perform well on the same sets of benchmarks.
@@
In order to understand whether the behavior of xMapleLCM-DIP was polynomial, we generated more challenging instances. Results can be seen in the left plot on the second row of Figure 6, with logarithmic scale on the $y$ axis. There is little doubt that the runtime of our solver ends up being exponential w.r.t. the size of the problem. However, we went one step further and studied how large the generated DRAT proofs were. In the same row, but on the right, we show the number of resolution steps in the proof. Despite having no theoretical support for that, it seems that although the solver takes exponential time in finding a proof, its size might be polynomial w.r.t. the problem size. That would indicate that our problem is having search heuristics that are not good enough to quickly find a short proof.
Finally, as we did in the previous section, we want to study the effect of the possible variants of DIP-based learning. For that purpose, we selected 5 challenging benchmarks of each type (grid, 4-regular, 6-regular) and evaluated the 7 DIP variants that we want to examine. Results can again be found in Table 1.
For Tseitin formulas, occ5 performs much better than the rest. Its only difference is that the minimum number of occurrences for a DIP to be introduced is smaller; this suggests that, in these instances, being more aggressive in the introduction of DIP extension variables provides a competitive advantage. On the other hand, glue solves no instance. We realized that DIP-learning was hardly ever applied, that is, requiring glue clauses was too restrictive here. Regarding act, whose performance is similarly poor, the main noticeable difference in its behavior is exhibiting a very large number of conflicts where DIP learning is applied, between 20 and 50%, which is around ten times more than the variants that performed best.
6.1.3 (X)Orified Random k-XOR Formulas
@@
One last family where DIP-based systems perform very well are random $k$-xor formulas, where orification or xorification [7] has been applied, i.e., replacing variables by (x)ors of fresh variables. We used CNFgen to obtain 4 sets of formulas: xor constraints of length 3 applying orification with 2 and 3 variables, and also with xorification of 2 and 3 variables. All formulas we consider are unsatisfiable, the number of variables before (x)orification is equal to the number of clauses, and we increase this number to get progressively more difficult formulas.
We studied the scaling properties of all systems by running them on increasingly larger benchmarks with a time limit of 300 seconds. Results can be seen in Figure 7, where the $y$ axis has logarithmic scale. Maybe surprisingly, CryptoMiniSat can only benefit from its preprocessing step in one of the families (2-xorified). Also, SBVA-CaDiCaL performs very well in two of them (2 and 3-orified). Again, both GlucoseER and our DIP-based implementations show the best average overall performance over all competitors.
Figure 7: Performance analysis of different solvers on (x)orified random $k$-xor formulas.
Finally, an analysis of the impact of the different DIP-learning variants can be found in Table 1. Our baseline implementation seems to perform the best, but there are several other configurations that perform very well. On the negative side, glue did not solve any such formula within the time limit. The identifying characteristic of that version is that the percentage of conflicts with DIP-learning is very low compared to the others (around 0.1%).
6.2 Performance Analysis on 2023 SAT Competition Benchmarks

### v2

@@ 5. Extension Variables from Dual Implication Points
5.1. An example with a grid Tseitin principle
Given a graph where every vertex has a charge, a number which is 00 or $1$, a Tseitin formula [Tse83] is created by considering one variable per edge, and adding one constraint per vertex $v$ expressing that the sum of the variables of all edges incident to $v$ modulo 2 is equal to its charge. The CNF version that we consider converts each xor constraint into clauses by simple enumeration of falsifying assignments. It is easy to see that the formula is unsatisfiable if and only if the sum of all charges is even. Here we consider as a graph the $3\times 3$-grid depicted in Figure 4 and only one vertex has charge 1, so the clauses are unsatisfiable.
@@ 5.1. An example with a grid Tseitin principle
Figure 4. An instance of the $3\times 3$-grid Tseitin principle. The nodes are assigned a polarity in $\{0,1\}$. The edges are labeled with variables.
@@
Figure 5. The complete conflict graph from the example; (a) at the first conflict and (b) at the second conflict. Boxed literals are decision literals. Implications at the top decision level are in bold.
It is interesting to relate the Figure 5(b) to Theorem 1 on DIPs. In this example, there is only one way to choose the two vertex-disjoint paths. Namely, to let $\pi_{a}$ and $\pi_{b}$ be the paths $e_{6},e_{8},e_{11},\overline{x},\overline{e_{7}},e_{10},\perp$ and $e_{6},e_{9},\overline{e_{12}},\perp$. The edge from $e_{6}$ to $\overline{e_{7}}$ bypasses $e_{8}$, $e_{11}$ and $\overline{x}$; and the edge from $e_{11}$ to $\overline{e_{12}}$ is a crossing separator. (The edge from $\overline{e_{7}}$ to $e_{9}$ is a “vacuous” crossing separator that does not actually remove any possible DIP pairs.)
@@
5.2. Extending CDCL with Dual Implication Points
@@
Filtering out bad-quality DIPs. Learning a DIP whenever we find one would be too prolific and overwhelm the solver. In order to determine whether the DIP chosen in the previous step had to be discarded, the first possibility we explored was again inspired by 1-UIP learning, where learning glue clauses, i.e. having LBD equal to $2$, is a desired situation. The first filtering mechanism we implemented discarded all DIPs that did not have a glue post-DIP clause. Another possibility we considered is to wait for a DIP to occur a certain number of times before using it in DIP-based learning. In our implementation, we tried different numbers as the minimum threshold to introduce a DIP. A third possibility is to use the activity-based heuristic of the literals in the DIP to assess its quality: DIPs whose literals have high decision-heuristic scores should be prioritized. In our implementation, we check whether the activity level of the current DIP is higher than the average activity level of the 20 most recently encountered DIPs; if so, the current DIP is a candidate for DIP-learning, otherwise it is discarded. We refer to these various techniques as glue, occ and act, respectively.
@@
It would make no sense to backtrack to level $\ell_{C}$ when $\ell_{C}<\ell_{D}$ since then neither $\lnot z$ nor $\lnot f$ would be propagated. However, another possible strategy would be to backtrack to the maximum of the decision levels $\ell_{C}$ and $\ell_{D}$. This would mean $\lnot z$ and $\lnot f$ are both propagated. The disadvantage of this when $\ell_{C}>\ell_{D}$ is that it would mean $\lnot z$ is asserted by the post-DIP clause at level $\ell_{C}$, whereas it could have been propagated at the previous decision level $\ell_{D}$. This breaks a usual invariant for CDCL solvers. This would only be possible in a solver that permits chronological backtracking [NR18, MB19]; xMapleLCM, however, does not support this.
The previous reasoning needs some clarification for the case when the extension variable $z$ with definition $z\leftrightarrow l_{1}\land l_{2}$ we want to use has already been introduced. If $z$ is undefined or defined at the current decision level nothing changes. We know it cannot be true at some previous level, because otherwise $l_{1}$ and $l_{2}$ would have been propagated at that level, and not in the current one as all literals belonging to a DIP do. If it is false at some previous level, then we have no guarantee that the pre or the post-DIP clause is asserting at any decision level. Fortunately, that rarely happens in practice. However, we can always perform standard 1UIP learning or try to apply DIP-based conflict analysis starting with the clause $z\lor\lnot l_{1}\lor\lnot l_{2}$ that we can guarantee is conflicting at the current decision level.
@@
Deleting extended variables. As it happens with lemma learning, where keeping too many lemmas slows down unit propagation, managing too many extended variables might also be counterproductive. Again following the analogy with learned lemmas, which are useful at some point of the search but might become inactive after a while, it is natural to think that extended variables follow the same behavior. All in all, it seems mandatory to consider the deletion of extended variables, which amounts to deletion all lemmas where they appear.
@@ 6. Experimental Evaluation
We have implemented the DIP-based clause learning schemes described in Section 5.2 on top of the xMapleLCM ER framework [^3] All sources used for this evaluation can be found in https://github.com/chjon/xMapleSAT/tree/main.. We started our experimental evaluation running a variant of DIP-based learning on benchmarks of the SAT Competition [BHI+23] and comparing it to MapleLCM [LLX+17], the CDCL SAT solver on which it is based. Even though there did not seem to be a systematic improvement on all benchmarks, a few families with important speedups were identified. We start this section by analyzing the impact of DIP-based learning on these families and then move to final considerations about the performance on the overall SAT competition benchmarks.
@@
For each benchmark family, we start by describing the problem they encode. After that, we evaluate the impact of the different techniques explained in Section 5.2. In particular, we first consider as a baseline a version that (i) finds DIPs in the middle of the conflict graph, (ii) only adds a DIP if it has occurred at least 20 times and (iii) always learns both the pre- and post-DIP learning clauses. Different variants are obtained by changing only one of the three previous design decisions at a time. Regarding the type of DIP used, we analyze the performance of closest, random, and heuristic , that is, the systems whose only difference w.r.t. baseline is that the type of DIP is changed. Regarding the criterion used to discard a DIP, we implemented the glue and the act configurations, but due to their poor performance we do not include these results. We present instead results about for the occ5 and occ50 configurations, which discard any DIP that has not occurred at least 5 or 50 times, respectively. Finally, we evaluated the performance of the 1-clause variant only learns the post-DIP clause.
After that, we report on the performance of a variety of state-of-the-art solvers, each with some distinguished characteristic: Kissat 4.0.2 [BFFH20] (an extremely efficient CDCL solver), CryptoMiniSat 5.12.1 [SNC09] (support for XOR reasoning), SBVA-CaDiCaL [HGH23] (introduction of new variables via SBVA and winner of the main track of the 2023 SAT Competition), GlucosER [AKS10] (extended-resolution based CDCL solver) and the best available configuration for xMapleLCM.
6.1. Tseitin Formulas
These formulas have already been described in Section 5.1. In this section, we consider three types of unsatisfiable formulas generated by CNFgen [LENV17]: (i) Tseitin formulas on the grid, where the graph is a rectangular grid and each vertex is connected to its 4 neighbors, except for the vertices on the boundaries of the grid which have fewer neighbors, (ii) 4-regular Tseitin formulas, where the graph is a $4$-regular random one, and (iii) 6-regular Tseitin formulas built on top of $6$-regular random graphs. The time limit was set to 600 seconds per instance.
@@ 6.1. Tseitin Formulas
Figure 6. Performance of different xMapleLCM configurations and state-of-the-art solvers on Tseitin formulas.
The first row of Figure 6 shows the effect of changing the type of DIP that is used on our baseline solver. A point $(x,y)$ indicates that this particular variant took $y$ seconds (note the logarithmic scale on the $y$ axis) to solve the corresponding Tseitin formula of size $x$. We observe that there is no big difference for regular Tseitin formulas. However, for formulas on the grid, the middle DIP is the one that scales best, being the only variant able to solve all instances within the time limit. In the second row of the same figure, we can see how learning both the pre- and the post-DIP clauses (baseline in the plots) is comparable to only learning post-DIP clause (1-clause). Finally, the third row shows that changing the minimum number of times a DIP has to occur before we introduce it has some impact. 50 seems to be too restrictive, whereas 20 (baseline) is slightly better than 5, which might end up adding too many DIPs.
The last row of Figure 6 compares the baseline configuration of xMapleLCM with state-of-the-art tools. The conclusion is clear: our DIP-based xMapleLCM solver is the only one that can process all instances within the time limit. GlucosER scales well on regular formulas, but not on formulas on the grid. Despite these formulas are in principle trivial for CryptoMiniSat, since the application of Gaussian elimination in a preprocessing step solves them, this does not seem to work for large formulas on the grid. Finally, Kissat and SBVA-CaDiCaL perform poorly, solving only very small instances.
@@
In order to understand whether the behavior of xMapleLCM was polynomial for these formulas, we generated more challenging Tseitin formulas on the grid. Results can be seen in the left plot of Figure 7, with logarithmic scale on the $y$ axis. There is little doubt that the runtime of our solver ends up being exponential w.r.t. the size of the problem. However, we went one step further and studied how large the generated DRAT proofs were. On the right plot, we show the number of resolution steps in the proof. Despite having no theoretical support for that, it seems that although the solver takes exponential time in finding a proof, its size might be polynomial w.r.t. the problem size. That would indicate that our problem is using search heuristics that are not good enough to quickly find a short proof.
Figure 7. Performance analysis of the baseline DIP-based solver on Tseitin formulas on the grid. The plot on the left represents running time (with logarithmic scale on $y$), whereas the one on the right shows the number of resolution steps in the generated proof.
6.2. Xorified Random k-XOR Formulas
@@
Another family where DIP-based systems perform very well are random $k$-xor formulas, where xorification [BS09] has been applied, i.e., replacing variables by xors of fresh variables. We used CNFgen to obtain 2 sets of formulas: xor constraints of length 3 applying xorification with 2 and 3 variables. All formulas we consider are unsatisfiable, the number of variables before xorification is equal to the number of clauses, and we increase this number to get progressively more difficult formulas. A time limit of 600 seconds was used.
The performance of the different xMapleLCM configurations can be seen in Figure 8. In the first row, we notice that there is not much difference if we change the DIP choice. A similar phenomenon can be observed in the second row, where the use of one or two learned DIP clauses does not have big impact on the performance. In the third row, we can see that changing the minimum number of occurrences of DIPs does affect the speed of the system: choosing 20 seems to give more consistent behavior.
Finally, in the last row of Figure 8 we notice that CryptoMiniSat can only benefit from its preprocessing step when xorification with 2 variables is applied. Our xMapleLCM solver is the one the performs best with GlucosER in second place. Finally, both SBVA-CaDiCaL and Kissat are only able to solve small instances.
Figure 8. Performance of different xMapleLCM configurations and state-of-the-arts solvers on XORified rand-kxor formulas.
6.3. Matching of Properly Intersecting Intervals
We are given a sequence of numbers $(a_{1},\ldots,a_{n})$, initially all set to zero, and a set of operations, each consisting of assigning a certain number to a contiguous subsequence of the $a_{i}$’s, defined by an interval. The goal is to perform each operation exactly once while maximizing the number of pairs of consecutive numbers $(a_{i},a_{i+1})$ that are different at the end of the process. Given some additional conditions, this can be formulated as maximum bipartite matching problem on a certain graph. By removing some edges from this graph, an unsatisfiable problem is generated. A more detailed description can be found in [BHI+23].
We downloaded the 23 instances submitted to the SAT Competition, using the Global Benchmark Database [IS19], and observed that no system without extended-resolution reasoning could solve any benchmark in a time limit of 2 hours. We used a much larger timeout than with the two previous families because for this family we are not able to construct increasingly larger benchmarks and see how the solvers scale.
Table 1 reports on our experimental evaluation. In the first block we can see that the random and heuristic DIP choices do significantly improve upon our baseline solver. The second block row shows how the behavior of the heuristic configuration can be further improved by increasing the threshold of minimum DIP occurrences to 50. On the contrary, a smaller threshold of 5 significantly degrades the performance, whereas only learning the post-DIP clause slightly slows it down. Finally, the third block compares our best xMapleLCM configuration with state-of-the-art solvers. It is remarkable that, apart from GlucosER, no system is able to process any instance within the given time limit. Hence, the use of extended resolution is crucial in this set of benchmarks. Our comparison shows that GlucosER and xMapleLCM solve the same number of instances, the former being only slightly faster.
Baseline Random Closest Heuristic
SOLVED 6 13 10 13
Average (secs.) 6011 4337 4818 4220
Heuristic 1-clause Occ-5 Occ-50
SOLVED 13 11 6 14
Average (secs.) 4220 4995 6057 3913
Occ-50 GlucosER Kissat CryptoMiniSat SBVA-CaDiCaL
SOLVED 14 14 0 0 0
Average (secs.) 3913 3622 7200 7200 7200
Table 1. Average running times in seconds of DIP-learning variants on properly intersecting intervals formulas. The best performing systems are in bold.
6.4. Performance Analysis on SAT Competition Benchmarks